the length of a rectangle is 8 cm. the perimeter is 18 cm. what is the area of the rectangle?

Answer by TimothyLamb(4379)

---
x = 8
2x + 2y = 18
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 8 cm
y = 1 cm
area = 8*1 = 8 sq.cm
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php